Twins graduate from college together and start their careers. Twin 1 invests $2000 at the end of each year for 10 years only (until age 32) in an account that earns 8%, compounded annually. Suppose that twin 2 waits until turning 40 to begin investing. How much must twin 2 put aside at the end of each year for the next 25 years in an account that earns 8% compounded annually in order to have the same amount as twin 1 at the end of these 25 years (when they turn 65)? (Round your answer to the nearest cent.) $
Answer #1

Answer:

Twin A/c balance after 31 years $28,973.12 =FV(8%,10,-2000,0,0)
Twin A/c balance after 65 years $396,645.88 =28973.12*(1+8%)^34
Twin 2 yearly deposit $5,425.63 =PMT(8%,25,0,-396645.88)